OpenStudy (anonymous):

f(ab)=f(a) + f(b) for all real numbers in the domain of f, f(x) equals which of these? (A)1/x (B) e^x (C) log(x)

OpenStudy (anonymous):

Try them and see?

OpenStudy (anonymous):

\[f(x) = \frac{1}{x}\implies \] \[f(ab) = \frac{1}{ab}\] \[f(a) + f(b) = \frac{1}{a} + \frac{1}{b} = \frac{a+b}{ab}\] \[\implies f(ab) \ne f(a)+f(b)\] Nope, not that one.

OpenStudy (anonymous):

Which one are you going to try next?

OpenStudy (anonymous):

would this process be right? ::: f(x)= e^x f(ab)= e^ab f(a)+f(b)= e^a +e^b which does not equal e^ab ?? yes?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

ohh~~ so it's the third choice! i see it now...i did log 5 + log 2 and then it came as one which is the same as log 5*2. i get it now. THANKS!!

OpenStudy (anonymous):

Yes, but cause the sum of logs is the log of the product.

OpenStudy (anonymous):

err because rather.

OpenStudy (anonymous):

\[log(a) + log(b) = log(a*b)\]